OpenStudy (anonymous):

solve by elimination method...-x+6y=5 and-3x+18y=15

OpenStudy (diana):

so -3(-x+6y=5) -3x+18y=15 ------------- 0=0 they are the same line

OpenStudy (anonymous):

ok so it is an empty set.